Determine the moment vector m about the origin, o, of a force f= 10i-j+2k passing through the point with position vector r=2i-3j+5k


1
Expert's answer
2022-05-09T16:17:33-0400
"\vec M=\vec {r}\times\vec{F}=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 2 & -3 & 5 \\
 10 & -1 & 2 \\
\end{vmatrix}"

"=\vec{i}\begin{vmatrix}
 -3 & 5 \\
 -1 & 2
\end{vmatrix}-\vec{j}\begin{vmatrix}
 2 & 5 \\
 10 & 2
\end{vmatrix}+\vec{k}\begin{vmatrix}
 2 & -3 \\
 10 & -1
\end{vmatrix}"

"=-\vec{i}+46\vec{j}+28\vec{k}"